Determine the equivalent total resistance for each of the following circuits below

7

5

2


1
Expert's answer
2022-03-14T13:06:31-0400

All resistance is series combination


"R_{eq}=R_1+R_2+R_3=7+5+2=14\Omega"

Parallel combination

"\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}"

"\frac{1}{R}=\frac{1}{7}+\frac{1}{5}+\frac{1}{2}"

"R=1.40\Omega"